5. A chemist has one solution that is 15% acid and another solution that is 25% acid. How much of the 15% solution is needed to make 100 L of a solution that is 18% acid?
a. 45 L b. 70 L c. 25 L d. 30 L

Found 2 solutions by Alan3354, stanbon:
Answer by Alan3354(69443)

A chemist has one solution that is 15% acid and another solution that is 25% acid. How much of the 15% solution is needed to make 100 L of a solution that is 18% acid?
------------
f = amount of 15%
t = amount of 25%
----------
f + t = 100 (total solution)
0.15f + 0.25t = 0.18*100 (total acid)
You can finish it.

Answer by stanbon(75887)

A chemist has one solution that is 15% acid and another solution that is 25% acid. How much of the 15% solution is needed to make 100 L of a solution that is 18% acid?
------------------------
Equation:
acid + acid = acid
0.15x + 0.25(100-x) = 0.18*100 L
---------------------------------------------
Multiply thru by 100 to get:
15x + 2500-25x = 1800
---
-10x = -700
x = 70 L (amt of 15% solution needed for the mixture)
---
100-x = 30L (amt of 25% solution needed for the mixture)
====================================
Cheers,
Stan H.


a. 45 L b. 70 L c. 25 L d. 30 L